OpenStudy (anonymous):

A basket contains 15 pieces of fruit: 7 apples, 5 oranges, and 3 bananas. Jonas takes a piece of fruit at random from the basket, and then Beth takes a piece at random. What is the probability that Jonas will get an orange and Beth will get an apple?

OpenStudy (experimentx):

5/15 for Jonas 7/14 for Beth

OpenStudy (carniel):

Jonas will be the 1st so he will have a probabillity of 7/15. Then Beth takes a Orange. That will me 5/14

OpenStudy (carniel):

*Sigh* Its the other way around my bad =D

OpenStudy (anonymous):

12 over 210

OpenStudy (carniel):

How did you get that?

OpenStudy (experimentx):

Oops sorry, 35/210 ??

OpenStudy (anonymous):

nvm its 1/6

OpenStudy (carniel):

It says whats the probabillity of them pull that certain fruit. It seems like you cross multiplied for both.

OpenStudy (experimentx):

yeah 35/210 = 1/6 => probability of Jonas x probability of Beth ... it's because they are connected by AND

OpenStudy (anonymous):

me no its 5/15 for jonas which is 1/3 and 7/14 for beth 1/2 in order to get the probablity compound i multiplied 1*1 and 3*2

OpenStudy (anonymous):

Since there are five oranges out of ten fruit the probability that John get an orange is 5/11.

OpenStudy (carniel):

Are you not looking for one person...

OpenStudy (anonymous):

@Rohangrr where u get 11 its 15

OpenStudy (anonymous):

sorry its 5/15 = 1/3